OpenStudy (anonymous):

The length of a rectangle is three times its width. If the area of the rectangle is 147 m^2 , find its perimeter.

OpenStudy (anonymous):



OpenStudy (anonymous):

so whats the perimeter?

OpenStudy (anonymous):

w = 7, L = 21 P = 2w + 2L= 14 + 42= 56